OpenStudy (anonymous):

integrate tan^7 (x/2) sec^2 (x/2) dx. pleaseee.

OpenStudy (anonymous):

d tan(x/2) =(1/2)sec^2 (x/2) integrate tan^7 (x/2) sec^2 (x/2) dx =integrate tan^7 (x/2) d 2tan(x/2) =2 integrate tan^7 (x/2) d tan(x/2) =2 (1/8)tan^8 (x/2) +C

OpenStudy (anonymous):

substitute\[u = \tan \left( \frac{x}{2} \right)\]\[du = \frac{1}{2}\sec^{2}\left( \frac{x}{2} \right)dx\]

OpenStudy (anonymous):

\(u=\tan(\frac{x}{2}), du=\frac{1}{2}\sec^2(\frac{x}{2})dx\) get \[\frac{1}{2}\int u^7du\] etc

OpenStudy (anonymous):

oops should be \(2\int u^7du\) since it is \(2du=\sec^2(\frac{x}{2})dx\)